OpenStudy (anonymous):

x^2=-x+6 (two answers)

OpenStudy (anonymous):

x^2 -6 + x =0

OpenStudy (anonymous):

A quadratic equation has two solutions and when they are in the equation, they are in the form of (x-a)(x-b) = 0. The solutions to this is x=a or x=b. Now consider this: x^2 + cx + d = 0 The solution to this is in the form shown above, which is (x-a)(x-b) = 0 The trick to get the solutions is this. when you multiply a and b, this should be equal to 'd' in this case and the sum of a and b should be equal to 'c'. Try this with your equation and you should get the answer.

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

My pleasure